40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Oct 26 21:46:33 2013, RANA, Update, IOO, Mode Cleaner Tune-UP MCtrend.pdfMC40days.png
    Reply  Mon Oct 28 14:12:48 2013, manasa, Update, IOO, Mode Cleaner relocked Screenshot-Untitled_Window-3.png
       Reply  Wed Oct 30 01:53:52 2013, Jenne, Update, IOO, Mode Cleaner relocked 
    Reply  Mon Oct 28 21:33:55 2013, RANA, Update, IOO, Mode Cleaner Tune-UP Untitled.png
       Reply  Thu Oct 31 15:34:28 2013, manasa, Update, IOO, MC not happy since yesterday MC_drift.png
       Reply  Thu Oct 31 20:05:48 2013, RANA, Update, IOO, Mode Cleaner Tune-UP drifty.png
          Reply  Fri Nov 1 09:45:32 2013, Steve, Update, IOO, beam dumps to be find razor_beam_dumps.jpg
          Reply  Mon Nov 4 11:37:12 2013, Steve, Update, IOO, PSL output shutter installed Shutterisback.jpg
             Reply  Mon Nov 4 12:59:43 2013, Jenne, Update, IOO, MC alignment not so good after PSL output shutter installed MCspots_4Nov2013.png
                Reply  Tue Nov 5 00:39:43 2013, manasa, Update, IOO, IFO alignment tuning MC_spot.png
                   Reply  Tue Nov 5 08:44:21 2013, Steve, Update, IOO, after last steering mirror mount swap after_Shutter_SMm_installed.pngafterlastSMmountswap.jpg
                   Reply  Tue Nov 5 17:12:34 2013, Steve, Update, LSC, centered qpds recentered.png
                      Reply  Wed Nov 6 08:33:53 2013, Steve, Update, IOO, Poiting changes of PSL 1dayTRENDofPOITING.png20dayTRENDioo.png
                         Reply  Thu Nov 7 08:57:20 2013, Steve, Update, IOO, PSL pointing monitoring PMC11072013.jpgPMCTqpdMon.png
                            Reply  Mon Nov 11 22:35:45 2013, RANA, Update, IOO, PSL pointing monitoring qpd.pdf
                               Reply  Tue Nov 12 23:48:23 2013, RANA, Update, IOO, PSL pointing monitoring 
                               Reply  Thu Nov 21 16:32:20 2013, Steve, Update, IOO, PMC needs a touch of love PMCplus.png
                               Reply  Fri Dec 6 15:57:41 2013, Steve, Update, IOO, beam dumps for PSL pointing monitoring  IOO_QPD_MONS.jpgPointingTrend23d.png
                                  Reply  Tue Dec 10 10:07:01 2013, Steve, Update, IOO, more beam traps InputWindowRefDump.jpgInpWindowRefDupm.jpg
Message ID: 9296     Entry time: Sat Oct 26 21:46:33 2013     Reply to this: 9303   9306
Author: RANA 
Type: Update 
Category: IOO 
Subject: Mode Cleaner Tune-UP 

 The MC had been unlocked for the last 4 hours and was crying out to me so I gave it some attention. Its happier now.

From the trend (AtM #1), I saw that the MC2 suspension has moved by ~10 microradians. Since the MC cavity divergence angle is lambda/(pi*w0) ~ 200 microradians, this isn't so much, but enough to cause it to lock on bad modes sometimes. Attackmint too shows that there's not much in monotonic drift over the last 40 nights.

I moved back MC2 to its old alignment with these commands:

ezcaservo -r C1:SUS-MC2_SUSPIT_INMON -s -1017 -g 0.0009 C1:SUS-MC2_PIT_COMM -t 300

ezcaservo -r C1:SUS-MC2_SUSYAW_INMON -s 490 -g 0.0009 C1:SUS-MC2_YAW_COMM -t 332

Then I went out to the table and aligned the beam into MC using the last two steering mirrors good enough so that the WFS coming on doesn't make the visibility any better. In this nominal state, I unlocked the MC and then aligned the reflected beam onto the center of the LSC PD as well as the WFS. The beam on the first WFS is a little small - next time someone wants to improve our Gouy phase telescope, we might try to make it bigger there. On the LSC PD, the beam was off-center by a few hundred microns.

Attachment 1: MCtrend.pdf  995 kB  Uploaded Sat Oct 26 22:47:18 2013  | Hide | Hide all
MCtrend.pdf
Attachment 2: MC40days.png  59 kB  Uploaded Sat Oct 26 22:56:50 2013  | Hide | Hide all
MC40days.png
ELOG V3.1.3-